OpenStudy (anonymous):

Please help!! Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4>

jimthompson5910 (jim_thompson5910):

Use the formula \[\Large \theta = \frac{u \cdot v}{|u||v|}\]

jimthompson5910 (jim_thompson5910):

|u| is the magnitude or length of vector u |v| is the magnitude or length of vector v

OpenStudy (anonymous):

but when i did this i got an answer that wasnt even one of the answer choices

jimthompson5910 (jim_thompson5910):

what did you get

jimthompson5910 (jim_thompson5910):

first tell me what the magnitudes of each vector are